# Inflationary power spectrum from the Lanczos algorithm

### v1

@@ I Introduction
[…]
@@
The quantum origin of curvature perturbation has attracted much attention [13, 14]. It motivates us to use alternative methods from quantum information to evaluate the curvature perturbation. One interesting direction is to implement the concept of complexity to study the curvature perturbations. There are various kinds of complexities in quantum information, such as computational complexity [15], Krylov complexity [46], and etc. The computational complexity in inflation was examined in [17, 18, 19, 20, 21, 22, 23]. Among of these, Ref. [24] has studied the Krylov complexity in cosmology with non-trivial sound speed. However, those literatures studied the complexity for the closed system. But our universe is an open system, therefore, one must use the open system approach to study the Krylov complexity in inflation [25, 26]. In [25, 26], the authors have successfully adopted the two-mode squeezed state to study the Krylov complexity in inflation. As Ref. [47] mentioned, the amplification from the quantum level to the classical level in inflation is a quantum squeezing process. Thus, it is quite natural to use the two-mode squeezed state to evaluate the quantum fluctuations in inflation. The two-mode squeezed state has been broadly applied to the pre-inflationary phase [27, 28, 29, 30, 31, 32].
@@
In this work, we will utilize the Bogoliubov transformation to establish a relationship between the two-mode squeezed state and the Bunch-Davies vacuum (BD). The Bogoliubov coefficients, denoted as $\alpha_{k}$ and $\beta_{k}$, are determined according to [33]. However, the wave function we employ is not the standard pure two-mode squeezed state; instead, we refer to it as the open two-mode squeezed state (OTMSS), which is derived from the second kind of Meixner polynomial (equivalent to the generalized Lanczos algorithm) and the symmetries of the Hamiltonian [26]. Following this, we then calculate the power spectrum of the OTMSS within the framework of the generalized Lanczos coefficient. And we find that the numerical results match those in the BD vacuum.
@@
This work is organized as follows. Sec. II will give some foundations of curvature perturbation under the framework of single-field inflation. Sec. III describes the generalized Lanczos algorithm in inflation. In Sec. IV, we will derive the so-called OTMSS. Sec. V will show the explicit formula of $\alpha_{k}$ and $\beta_{k}$. Then, Sec. VI will obtain the numerics of $\phi_{k}$ and $r_{k}$. Based on this, Sec. VII will derive the power spectrum of OTMSS. Finally, the conclusion and discussion are drawn in Sec. VIII.
@@ II Some foundation of inflationary perturbation
In this work, we will follow Ref. [34] to give some foundation for the cosmological perturbation. First, it is the spatially conformal flat Friedmann-Lemaitre-Robertson-Walker (FLRW) metric,
@@
where $\psi(\eta,\vec{x})$ is the perturbation of metric and $a(\eta)$ is the scale factor. Meanwhile, we can also define the perturbation of the inflaton as $\phi(x_{\mu})=\phi_{0}(\eta)+\delta\phi(x_{\mu})$, where $\phi_{0}(\eta)$ is the background field of inflaton and the $\delta\phi(x_{\mu})$ can be dubbed as the perturbation of inflaton. Being armed with these perturbations, we can obtain the generically quadratic action as follows,
[…]
where the Hubble rate $H=\dot{a}/a$ (denoting the derivation with the physical time), the conserved curvature perturbations is $\mathcal{R}=\psi+\frac{H}{\phi_{0}}\delta\phi$ at horizon crossing $(k=aH)$ and the Hubble flow expansion $\epsilon=-\dot{H}/H^{2}$, $z=\sqrt{2\epsilon}a$. Furthermore, we rewrite action (2) in terms of the Mukhanov-Sasaki variable $v=z\mathcal{R}$,
[…]
@@
However, defining a privileged quantum vacuum state in curved spacetime is fundamentally problematic, as established by [35]. In this context, the BD vacuum emerges as a particular state that asymptotically approaches the Minkowski vacuum at short distances [36]. The complete vacuum characterization requires consideration of both the Poincaré patch of de Sitter spacetime (PdS) and Hadamard regularity conditions. Following standard cosmological practice, we maintain the BD vacuum assumption under PdS invariance and Hadamard conditions, consistent with foundational work [37, 38, 33].
@@ III Lanczos algorithm in inflation
Recent studies [25, 39, 40, 26, 41] have demonstrated that the early universe behaved as a dissipative system, owing to both the violation of energy conservation [42] and the particle production during preheating [43, 44]. Motivated by these findings, we employ the Lanczos algorithm for open quantum systems to evaluate the quadratic action (3). This approach aligns with the generalized Lanczos algorithm framework developed in [45]. Following our previous conventions [25], we express the generalized Lanczos algorithm as
@@
where the coefficients $c_{n}$ encode the open quantum system dynamics for a given Hamiltonian, while $b_{n}$ (the Lanczos coefficients) characterize the chaotic properties of the dynamical system. The orthogonal basis $\mathcal{O}_{n}=|n,n\rangle_{-\vec{k},\vec{k}}$ spans the Krylov space, where it denotes the space of the two-mode squeezed state, and $\mathcal{L}$ denotes the Liouvillian superoperator – which can be represented in terms of creation and annihilation operators analogous to a Hamiltonian operator.
Following our previous work [25, 39, 40], we identify that:
  • The $c_{n}$ terms in (6) correspond to open system effects
  • The $b_{n}$ terms represent closed system dynamics
@@
    Isolate environmental decoherence effects (open system)
@@
For comprehensive discussions of the Lanczos algorithm implementation and Krylov complexity in this context, we refer readers to [46].
[…]
@@
where $\hat{c}_{-\vec{k}}^{\dagger}$ and $\hat{c}_{\vec{k}}$ denote the creation and annihilation operators, respectively, $u_{k}$ and $v_{k}$ are the function of conformal time $\eta$. Here, we keep two modes $(k,-k)$ since it is invariant under changing from $k$ to $-k$ in action (3) like in [36]. Therefore, the Hamiltonian (8) will become as follows
$\begin{split}\hat{H}=\int{d^{3}k}\hat{H}_{k}=&\int{d^{3}k}\bigg{[}k\hat{c}_{-\vec{k}}^{\dagger}\hat{c}_{-\vec{k}}+k\hat{c}_{\vec{k}}\hat{c}_{\vec{k}}^{\dagger}\\ &+i\frac{z{}^{\prime}}{z}\hat{c}_{\vec{k}}^{\dagger}\hat{c}_{-\vec{k}}^{\dagger}-i\frac{z{}^{\prime}}{z}\hat{c}_{\vec{k}}\hat{c}_{-\vec{k}}\bigg{]}.\end{split}$ (11)
where have chosen appropriate normalizations for $v_{\vec{k}}$ and $u_{\vec{k}}$ and this Hamiltonian is of the Weyl-group structure [57]. Then, we can use this Hamiltonian (11) to act on the two-mode orthogonal basis $\mathcal{O}_{n}=|n,n\rangle_{-\vec{k},\vec{k}}$, one can obtain
$\begin{split}\hat{H}|\mathcal{O}_{n})=(2n+1)k|\mathcal{O}_{n})+i\frac{z{}^{\prime}}{z}(n+1)|\mathcal{O}_{n+1})\\ -in\big{(}\frac{z{}^{\prime}}{z}\big{)}|\mathcal{O}_{n-1}).\end{split}$ (12)
where $b_{n}^{2}=n^{2}(\frac{z^{\prime}}{z})^{2}$, $c_{n}=i(2n+1)k$ and the prime symbol indicates differentiation with respect to conformal time. The formula of Lanczos coefficient clearly indicates that the early universe is a maximally chaotic, infinity dynamical system as discovered in [25, 39, 40, 26]. For convenience, we will define
@@
where we have followed the notation of [25] and restored the unit of Planck mass, $\mu_{2}$ plays a role of dissipative coefficient and thus $b_{n}^{2}=|1-\mu_{1}^{2}|$ and $c_{n}=i(2n+1)\mu_{2}$. As previously mentioned, the $c_{n}$ part corresponds to the open system part of Hamiltonian (11), and the $b_{n}$ matches the closed system part of Hamiltonian (11). Thus, we could split the Hamiltonian into the open part and closed part as follows,
@@
As Ref. [47] pointed out, the amplification from the quantum level to the classical level in inflation is a quantum squeezing process. Therefore, it is natural to use the two-mode squeezed state to evaluate the quantum fluctuations in inflation. However, in this work, we will implement a more generic function to replace this wave function. First and foremost, the so-called OTMSS via Ref. [25] is
@@
where $|n,n\rangle_{-\vec{k},\vec{k}}$ is the Krylov basis in this work, $\eta$ is the conformal time, $r_{k}$ and $\phi_{k}$ are two squeezing parameters, $\mu_{2}$ and $|1-\mu_{1}^{2}|$ as defined in Eq. (13). Here, we will introduce more information on wave function (15), in which this wave function will describe the quantum fluctuations in inflation. First, this wave function could represent the dissipative effects of one dynamical system denoted by $\mu_{2}$, and it will nicely recover the two-mode squeezed state in the weak dissipative approximation $(\mu_{2}\ll 1)$,
@@
Thus, we can see that the pure two-mode squeezed state is a wave function that corresponds to a closed system. In later illustrations, we will confirm this point. Secondly, the derivation of the OTMSS (15) only relies on the second kind of Meixner polynomial [48],
@@
where we define $\tilde{c}_{n}=-ic_{n}$, and $x$ denotew the Hamiltonian. The sequence begins with the initial conditions $P_{0}(x)=1$ and $P_{1}(x)=x-c_{0}$. The polynomial $P_{n}(x)=\det(x-\mathcal{L}_{n})$ corresponds to an open system, and $\mathcal{L}_{n}$ represents the Liouvillian superoperator for the $n$-th quantum state. By introducing the natural orthonormal basis $\{e_{n}\}$. We could represent $|P_{n}(x))=\Bigl{(}\prod_{k=1}^{n}b_{k}\Bigr{)}|\mathcal{O}_{n}),\ \ \mbox{and}\ \ \ |x^{n})=\mathcal{L}^{n}|\mathcal{O}).$ Combining Eq. (17), there is an explicit relation as follows,
@@
where $e_{n}=\mathcal{O}_{n}$ in our case and then it is equivalent to the generalized Lanczos algorithm (6). Through this simple illustration, one could explicitly see that the OTMSS only relies on the second kind of Meixner polynomial, describing the dissipative effects of open system (the early universe). Therefore, we could see that the OTMSS (15) is more generic compared with the pure two-mode squeezed state. Finally, we need to emphasize that the OTMSS is model-independent since the information of various models is encoded in $\mu_{2}$ and $|1-\mu_{1}^{2}|$.

### v2

@@ Abstract
The inflationary paradigm Guth:1980zm ; Starobinsky:1980te ; Linde:1981mu addresses several fundamental issues in standard Big Bang cosmology, including the flatness problem, the horizon problem, and the magnetic monopole problem. Quantum fluctuations in the vacuum state during quasi-de Sitter expansion predict the power spectrum of primordial curvature perturbations Mukhanov:1990me ; Martin:2024qnn , which is nearly scale-invariant as constrained by observations of the cosmic microwave background (CMB) Planck:2018jri . The advent of high-precision CMB measurements has initiated an era of precision cosmology. Consequently, current and forthcoming observations are capable of probing primordial curvature perturbations beyond the predictions of standard inflationary models. CORE:2016ymi ; CMB-S4:2020lpa ; LiteBIRD:2022cnt ; Euclid:2023shr ; Mergulhao:2023ukp ; Antony:2024vrx .
@@
The quantum origin of curvature perturbations has attracted significant attention Balitsky:2014epa ; Dalianis:2024 . This motivates the application of alternative methods from quantum information theory to evaluate curvature perturbations. One promising direction involves utilizing the concept of complexity to study these perturbations. Several types of complexity exist in quantum information, including computational complexity Papadimitriou:2003 , Krylov complexity Parker:2018yvk , among others. Computational complexity in the context of inflation has been investigated in Choudhury:2020hil ; Bhargava:2020fhl ; Bhattacharyya:2020rpy ; Adhikari:2021ked ; Li:2023ekd ; Liu:2021nzx ; Li:2021kfq . Among these, Ref. Adhikari:2022oxr studied Krylov complexity in cosmology with non-trivial sound speed. However, these works primarily address closed quantum systems. Since our universe is fundamentally an open system, it is essential to adopt an open quantum systems approach to study Krylov complexity during inflation Li:2024kfm ; Zhai:2024odw . In Li:2024kfm ; Zhai:2024odw , the authors successfully employed the two-mode squeezed state formalism to analyze Krylov complexity in inflation. As emphasized in Grishchuk:1990bj , the amplification of quantum fluctuations to classical scales during inflation is effectively a quantum squeezing process. Therefore, the two-mode squeezed state provides a natural framework for describing quantum fluctuations in this context. This approach has also been extensively applied to study the pre-inflationary phase Kowalczyk:2024ech ; Vilenkin:1982wt ; Starobinsky:2001kn ; Gessey-Jones:2021yky ; Handley:2016ods ; Contaldi:2003zv .
@@
In this work, we employ the Bogoliubov transformation to relate the two-mode squeezed state to the Bunch-Davies (BD) vacuum. The Bogoliubov coefficients, $\alpha_{k}$ and $\beta_{k}$, are determined according to Bunch:1978yq . The wave function used here is not the conventional pure two-mode squeezed state, but rather the open two-mode squeezed state (OTMSS), constructed via the second kind of Meixner polynomial, which is equivalent to the generalized Lanczos algorithm along with the symmetries of the Hamiltonian Zhai:2024odw . Within this framework, we compute the power spectrum of the OTMSS using the generalized Lanczos coefficients. Our numerical results show agreement with those obtained in the BD vacuum.
@@
This work is organized as follows. Sec. II will give some foundations of curvature perturbation under the framework of single-field inflation. Sec. III describes the generalized Lanczos algorithm in inflation. In Sec. IV, we will review the so-called OTMSS. Sec. V will show the explicit formula of $\alpha_{k}$ and $\beta_{k}$. Then, Sec. VI will obtain the numerics of $\phi_{k}$ and $r_{k}$. Based on this, Sec. VII will derive the power spectrum of OTMSS. Finally, the conclusion and discussion will be given in Sec. VIII.
@@ II Some foundation of inflationary perturbation
In this work, we will follow Ref. Baumann:2009ds to give some foundation for the cosmological perturbation. First, it is the spatially conformal flat Friedmann-Lemaitre-Robertson-Walker (FLRW) metric,
@@
where $\psi(\eta,\vec{x})$ denotes the metric perturbation and $a(\eta)$ is the scale factor. We also define the perturbation of the inflaton field as $\phi(x_{\mu})=\phi_{0}(\eta)+\delta\phi(x_{\mu})$, where $\phi_{0}(\eta)$ is the homogeneous background inflaton and $\delta\phi(x_{\mu})$ represents the inflaton perturbation. Using these definitions, the generic quadratic action can be derived as follows:
$S=\frac{1}{2}\int dtd^{3}xa^{3}\frac{\dot{\phi}}{H^{2}}\bigl{[}\dot{\mathcal{R}}-\frac{1}{a}(\partial_{i}\mathcal{R})^{2}\bigr{]},$ (2)
where the Hubble parameter is defined as $H=\dot{a}/a$ (denoting the derivation with the physical time), the conserved curvature perturbations is $\mathcal{R}=\psi+\frac{H}{\phi_{0}}\delta\phi$ at horizon crossing $(k=aH)$ and the slow-roll parameter is $\epsilon=-\dot{H}/H^{2}$, which leads to $z=\sqrt{2\epsilon}a$. Furthermore, the action (2) can be rewritten in terms of the Mukhanov-Sasaki variable $v=z\mathcal{R}$,
$S=\frac{1}{2}\int d\eta d^{3}x[v^{\prime 2}-(\partial_{i}v)^{2}+(\frac{z^{\prime}}{z})^{2}v^{2}-2v^{\prime}v\frac{z^{\prime}}{z}].$ (3)
@@
However, the definition of a privileged quantum vacuum state in curved spacetime is fundamentally ambiguous, as emphasized in Unruh:1976db . In this setting, the BD vacuum arises as a specific state that approximates the Minkowski vacuum at short distances Agullo:2022ttg . A complete characterization of the vacuum requires the framework of the Poincaré patch of de Sitter spacetime (PdS) together with Hadamard regularity conditions. In accordance with standard cosmological practice, we adopt the BD vacuum under the assumptions of PdS invariance and Hadamard regularity, following the foundational treatments in Chernikov:1968zm ; Tagirov:1972vv ; Bunch:1978yq .
@@ III Lanczos algorithm in inflation
Recent studies Li:2024kfm ; Li:2024iji ; Li:2024ljz ; Zhai:2024odw ; Shandera:2017qkg have demonstrated that the early universe behaved as a dissipative system, owing to both the violation of energy conservation Cheung:2007st and the particle production during preheating Kofman:1997yn ; Kofman:1994rk . Motivated by these findings, we employ the Lanczos algorithm for open quantum systems to evaluate the quadratic action (3). There are multiple approaches to Krylov complexity in open quantum systems. In this work, we employ the generalization of the Lanczos algorithm to the Arnoldi iteration via the Heissenberg form Bhattacharjee:2022lzy ; Bhattacharya:2022gbz ; Liu:2022god ; Nizami:2023dkf ; Nandy:2024evd , as this framework has proven effective for generating the OTMSS in our previous studies Li:2024kfm ; Li:2024iji ; Li:2024ljz ; Zhai:2024odw , particularly building on the methodology established in Bhattacharya:2022gbz . We begin by introducing the general Lindblad master equation Lindblad:1975ef :
$\dot{\rho}=-i[H,\rho]+\sum_{k}[L_{k}\rho L_{k}^{\dagger}-\frac{1}{2}\{L_{k}^{\dagger}L_{k},\rho\}],$ (6)
where $\rho$ is the density matrix, $H$ is the Hamiltonian, and $L_{k}^{\dagger}$ denotes the jump operator (which corresponds to the annihilation operator in our context). Following Bhattacharjee:2022lzy , we assume the open system dynamics occurs in an infinite temperature environment-modeled by the maximally mixed state. This serves as a reasonable approximation since the temperature of the inflationary universe is exceedingly high. Consequently, we obtain
$\sum_{k}[L_{k}\rho_{\infty}L_{k}^{\dagger}-\frac{1}{2}\{L_{k}^{\dagger}L_{k},\rho_{\infty}\}]=0,$ (7)
where $\rho_{\infty}$ denotes the infinite temperature-state. Under this assumption, the operator dyamics is governed by
$\mathcal{\hat{O}}(t)=\exp(i\mathcal{L}t)\mathcal{\hat{O}}$, (8)
where the Heisenberg-picture Lindbladian $\mathcal{L}$ is given by Bhattacharya:2022gbz ,
$\mathcal{L}=\mathcal{L}_{H}+\mathcal{L}_{D},~~\mathcal{L}_{H}\mathcal{\hat{O}}=[H,\mathcal{\hat{O}}],~~\mathcal{L}_{D}=\sum_{k}[L_{k}\mathcal{\hat{O}}L_{k}^{\dagger}-\frac{1}{2}\{L_{k}^{\dagger}L_{k},\mathcal{\hat{O}}\}$] (9)
where $\mathcal{L}_{H}$ governs the Hermitian dynamics, $\mathcal{L}_{D}$ represents the dissipative part, and $\mathcal{L}$ acts as a superoperator on the operator space. The inner product can be defined as $(A|B):=\mathrm{Tr}[\rho_{\infty}A^{\dagger}B]$, which explicitly connects the operator space structure to the Krylov basis.
Secondly, an appropriate basis is required to compute the aforementioned physical quantities. Following Bhattacharya:2022gbz , we adopt an orthonormal basis $\{\mathcal{O}_{0},\ldots,\mathcal{O}_{n},\ldots\}$ generated by the open-system Lindbladian according to the following procedure:
$\rm span(\mathcal{O}_{0},...,\mathcal{O}_{n})=span(\mathcal{V}_{0},...,\mathcal{V}_{n}),$ (10)
where $\mathcal{V}_{n}$ is the original Krylov basis, and this basis can be generated by the Arnoldi iteration. Here, we will recall the main idea which is started with an initial normalized vector $\mathcal{O}_{0}\propto\mathcal{O}$. Then, we could construct the subsequent Krylov basis vectors by repeatedly applying the operator and orthonormalizing at each step.
$|\mathcal{U}_{k})=\mathcal{L}|\mathcal{O}_{k-1}).$ (11)
From $j=0$ to n1n-1, one can perfom the following iterations: $1.~h_{j,k-1}=(\mathcal{O}_{j}|\mathcal{U}_{k})$, $2.~|\tilde{\mathcal{U}}_{k})=|\mathcal{U}_{k})-\sum_{j=0}^{k-1}h_{j,k-1}|\mathcal{O}_{j})$, $3.~h_{k,k-1}=\sqrt{(\tilde{U}_{k}|\tilde{\mathcal{U}}_{k})}$, it will stop as if $h_{k,k-1}=0$, otherwise, we could define
$|\mathcal{O}_{k})=\frac{|\tilde{\mathcal{U}}_{k})}{h_{k,k-1}}.$ (12)
Consequently, the Lindbladian can transform into o an upper Heissenberg form in the span Krylov basis (also called the Arnoldi basis)
$\mathcal{L}=\begin{pmatrix}h_{0,0}&h_{0,1}&h_{0,2}&\cdots&\cdots&h_{0,n}\\ h_{1,0}&h_{1,1}&h_{1,2}&\cdots&\cdots&h_{1,n}\\ 0&h_{2,1}&h_{2,2}&h_{2,3}&\cdots&\cdots\\ \vdots&0&h_{3,2}&\cdots&\cdots&\cdots\\ 0&\vdots&0&\vdots&\ddots&h_{n-1,n}\\ 0&0&\vdots&0&h_{n,n-1}&h_{n,n}\end{pmatrix}$ (13)
where $h_{m,n}=(\mathcal{O}_{m}|\mathcal{L}|\mathcal{O}_{n})$. When $\mathcal{L}$ is a Hermitian superoperator, the Arnoldi iteration will nicely recover into the Lanczos algorithm. Following our previous conventions Li:2024kfm , we express the generalized Lanczos algorithm as
@@
where the coefficients $c_{n}$ capture the open quantum system dynamics for a given Hamiltonian (corresponding to the diagonal elements $h_{n,n}$ in (13)), while the Lanczos coefficients $b_{n}$ (associated with the off-diagonal elements $h_{n,n-1}$ in (13)) characterize the chaotic behavior of the dynamical system. The orthogonal basis $\mathcal{O}_{n}=|n,n\rangle_{-\vec{k},\vec{k}}$ spans the Krylov space defined by Eq. (10). The Liouvillian superoperator $\mathcal{L}$, which admits a representation in terms of creation and annihilation operators equivalent to that of a Hamiltonian operator, which governs the time evolution.
@@
    Isolate the part of open system
@@
For comprehensive discussions of the Lanczos algorithm implementation and Krylov complexity in this context, we refer readers to Parker:2018yvk .
To connect with the generalized Lanczos algorithm, we need to represent the action (3) into the Hamiltonian in terms of creation and annihilation operators, where the definition of Hamiltonian is $H=\int d^{3}xd\eta(\pi v^{\prime}-L)$ with the conjugate momentum as
@@
where $\hat{c}_{-\vec{k}}^{\dagger}$ and $\hat{c}_{\vec{k}}$ represent the creation and annihilation operators, respectively, and $u_{k}(\eta)$, $v_{k}(\eta)$ are functions of conformal time $\eta$. We retain both modes $(\vec{k},-\vec{k})$ due to the invariance of the action (3) under $\vec{k}\rightarrow-\vec{k}$, following the treatment in Agullo:2022ttg . Accordingly, the Hamiltonian (16) takes the following form:
$\begin{split}\hat{H}=\int{d^{3}k}\hat{H}_{k}=\int{d^{3}k}\bigg{[}k\hat{c}_{-\vec{k}}^{\dagger}\hat{c}_{-\vec{k}}+k\hat{c}_{\vec{k}}\hat{c}_{\vec{k}}^{\dagger}+i\frac{z{}^{\prime}}{z}\hat{c}_{\vec{k}}^{\dagger}\hat{c}_{-\vec{k}}^{\dagger}-i\frac{z{}^{\prime}}{z}\hat{c}_{\vec{k}}\hat{c}_{-\vec{k}}\bigg{]}.\end{split}$ (19)
where we have chosen appropriate normalizations for $v_{\vec{k}}$ and $u_{\vec{k}}$ and this Hamiltonian is of the $SL(2,R)$ structure Caputa:2021sib . We then act with this Hamiltonian (19) on the two-mode orthogonal basis $\mathcal{O}_{n}=|n,n\rangle_{-\vec{k},\vec{k}}$, yielding:
$\begin{split}\hat{H}|\mathcal{O}_{n})=(2n+1)k|\mathcal{O}_{n})+i\frac{z{}^{\prime}}{z}(n+1)|\mathcal{O}_{n+1})-in\big{(}\frac{z{}^{\prime}}{z}\big{)}|\mathcal{O}_{n-1}).\end{split}$ (20)
where
$b_{n}^{2}=n^{2}(\frac{z^{\prime}}{z})^{2},~~~~c_{n}=i(2n+1)k,$ (21)
The prime symbol denotes differentiation with respect to conformal time. As shown in Parker:2018yvk , the dynamical system becomes maximally chaotic and infinite when $b_{n}\propto n$. Hence, the expression for the Lanczos coefficient in (21) clearly demonstrates that the early universe constitutes a maximally chaotic, infinite dynamical system, as established in Li:2024kfm ; Li:2024iji ; Li:2024ljz ; Zhai:2024odw . For convenience, we define
@@
where we follow the notation of Li:2024kfm and restore the Planck mass units. Here, $\mu_{2}$ acts as a dissipative coefficient, leading to the expressions $b_{n}^{2}=n^{2}|1-\mu_{1}^{2}|$ and $c_{n}=i(2n+1)\mu_{2}$. As noted earlier, the $c_{n}$ term corresponds to the open system contribution of the Hamiltonian (19), while the $b_{n}$ term aligns with the closed system part. We can therefore decompose the Hamiltonian into open and closed parts as follows:
@@
Here, we propose a new perspective on the generalized Lanczos algorithm in this work. This algorithm explicitly highlights the open system contribution to the quadratic perturbation action in (19), an aspect often overlooked in conventional treatments, where the dissipative coefficient is defined in direct relation to various Hamiltonian components.
@@
As noted in Ref. Grishchuk:1990bj , the amplification from quantum to classical scales during inflation is fundamentally a quantum squeezing process. It is therefore natural to employ the two-mode squeezed state to describe quantum fluctuations in this context. In this work, however, we adopt a more general wave function. Specifically, we use the OTMSS, as introduced in Li:2024kfm , defined as follows:
@@
where $|n,n\rangle_{-\vec{k},\vec{k}}$ denotes the Arnoldi basis used in this work, $\eta$ is the conformal time, $r_{k}$ and $\phi_{k}$ are the squeezing parameters, and $\mu_{2}$ and $|1-\mu_{1}^{2}|$ are as defined in Eq. (22). We now provide further details on the wave function given in Eq. (24), which describes quantum fluctuations during inflation. This wave function incorporates dissipative effects through the parameter $\mu_{2}$, and reduces to the standard two-mode squeezed state in the weak dissipative limit ($\mu_{2}\ll 1$),
@@
It follows that the pure two-mode squeezed state corresponds to a wave function describing a closed quantum system, a point we will reaffirm in subsequent discussions. Secondly, the OTMSS given in Eq. (24) is constructed based on the second kind of Meixner polynomial, Hetyei_2009 ,
@@
where we define $\tilde{c}_{n}=-ic_{n}$, and $x$ denote the Hamiltonian. The sequence begins with the initial conditions $P_{0}(x)=1$ and $P_{1}(x)=x-c_{0}$. The polynomial $P_{n}(x)=\det(x-\mathcal{L}_{n})$ corresponds to an open system, and $\mathcal{L}_{n}$ represents the Liouvillian superoperator for the $n$-th quantum state. By introducing the natural orthonormal basis $\{e_{n}\}$. We could represent $|P_{n}(x))=\Bigl{(}\prod_{k=1}^{n}b_{k}\Bigr{)}|\mathcal{O}_{n}),\ \ \mbox{and}\ \ \ |x^{n})=\mathcal{L}^{n}|\mathcal{O}).$ Combining Eq. (26), there is an explicit relation as follows,
@@
where $e_{n}=\mathcal{O}_{n}$ in our construction, which is equivalent to the generalized Lanczos algorithm given in Eq. (14). From this straightforward representation, it is evident that the OTMSS is constructed entirely from the second kind of Meixner polynomial, capturing dissipative effects inherent to open quantum systems such as the early universe. This demonstrates that the OTMSS in Eq. (24) is more general than the pure two-mode squeezed state. Moreover, the OTMSS is model-independent, as the parameters $\mu_{2}$ and $|1-\mu_{1}^{2}|$ encode specific physical information from different cosmological scenarios.
In this section, we have already checked that the OTMSS is more generic and complete compared to the two-mode squeezed state. Before presenting the conceptual advantage of the OTMSS, we need to provide additional explanation for the two-mode squeezed state. To proceed, we will define a generalized displacement operator as follows:
$\begin{split}D(\eta)&\ =e^{iHt}=\exp(ikt(c_{k}c_{k}^{\dagger}+c_{-k}^{\dagger}c_{-k})-\frac{z^{\prime}}{z}t(c_{-k}^{\dagger}c_{k}^{\dagger}-c_{-k}c_{k}))\\ &\ =\exp(-\eta c_{k}^{\dagger}c_{-k}^{\dagger})\exp(\bar{\eta}c_{-k}c_{k})\exp(-\alpha(c_{k}c_{k}^{\dagger}+c_{-k}^{\dagger}c_{-k})),\end{split}$ (28)
where $t$ is just a parameter, $\hat{H}$ is our Hamiltonian operator (19) and we also define $\eta=\frac{z^{\prime}}{z}t-i\frac{z^{\prime}}{z}kt^{2}$ and $\alpha=\frac{1}{2}(\frac{z^{\prime}}{z}t-\frac{z^{\prime}}{z}kt^{2})^{2}-ikt$. After some tedious calculations, we could obtain
$\begin{split}|\mathcal{O}(t))&\ =\exp(-\eta c_{k}^{\dagger}c_{-k}^{\dagger})\exp(\bar{\eta}c_{-k}c_{k})\exp(-\alpha(c_{k}c_{k}^{\dagger}+c_{-k}^{\dagger}c_{-k}))|0,0\rangle_{k,-k}\\ &\ =b\sum_{n=0}^{\infty}\frac{(-1)^{n}}{n!}e^{i2n\phi}\tanh^{n}r(c_{-k}^{\dagger}c_{k}^{\dagger})^{n}|0,0\rangle_{k,-k}\\ &\ =b\sum_{n=0}^{\infty}(-1)^{n}e^{i2n\phi}\tanh^{n}r|n,n\rangle_{k,-k},\end{split}$ (29)
where we have set $\eta=e^{2i\phi}\tanh r$ in complex coordinates, $r=|\eta|$ and $b$ is the normalization factor that can be derived as
$1=(\mathcal{O}(t)|\mathcal{O}(t))=|b|^{2}\sum_{n=0}^{\infty}\tanh^{2n}r=|b|^{2}\frac{1}{1-\tanh^{2}r}=|b|^{2}\cosh^{2}r,$ (30)
we obtain $b=\frac{1}{\cosh r}$. This result demonstrates that only the $H_{\text{closed}}$ component constructs the conventional two-mode squeezed state. Consequently, the OTMSS, derived from the second kind of Meixner polynomial, represents a generalization of the standard two-mode squeezed state. To our knowledge, this is the first time the OTMSS has been applied to compute the inflationary power spectrum. This generalization naturally incorporates dissipative effects, as described by the $H_{\text{open}}$ term in Eq. (23).